OpenStudy (anonymous):

Simplify: 3 4/5 + 8 4/5 Please ???

OpenStudy (anonymous):

3 + 8 and 4/5 + 4/5

OpenStudy (anonymous):

11 8/5

OpenStudy (anonymous):

12 3/5

OpenStudy (jiteshmeghwal9):

change this improper fractions into mixed fractions\[3 \frac{4}{5}=\frac{5 \times 3 +4 }{5}=\]\[8 \frac{4}{5}=\frac{5 \times 8 +4}{5}\]

OpenStudy (jiteshmeghwal9):

\[\frac{19}{5}+\frac{44}{5}\]now take this denominator common\[\frac{19+44}{5}=\frac{63}{5}\]